OpenStudy (otonogold):

Find the area of a parallelogram if the area is represented by x2 - 64, the length of a side is represented by x - 12, and the length of the altitude to that side is represented by x + 13.

OpenStudy (agent0smith):

Oh you're meant to solve it i guess, area is base*height: Area = (x - 12)(x + 13) = x^2 - 64 So solve (x - 12)(x + 13) = x^2 - 64 for x, by expanding out and simplifying etc, then plug x into Area = x^2-64